OpenStudy (anonymous):

how do you find out if lines are parallel, perpendicular, or neither? ex: 9x+3y=6 3x+9y=6

OpenStudy (anonymous):

convert the line inti the form Y=mX+c m=gradient if the gradient of both line is same then they are parallel if the gradient of L1 x the gradient of L2=-1 they are perpendicular

OpenStudy (anonymous):

gradient=slope i knw that sir

OpenStudy (anonymous):

he meant that for me lol. so i have to change both the lines into y = mx + b?

OpenStudy (anonymous):

not trying to be mean but considering the question saying gradient is confusing since you don't learn that till college

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

convert to Y=mX+c